# T3: Benchmarking Sycophancy and Skepticism in Causal Judgment

### v1

@@ 1.1 Relation to Prior Work
Two research threads inform our approach: formal causal assessment (Spirtes et al., 2000; Schölkopf et al., 2021) and the study of alignment pathologies (Bai et al., 2022).
@@
Several benchmarks evaluate LLMs on Pearl’s hierarchy. CLadder (Jin et al., 2023) generates queries from causal graphs with oracle labels across rungs, offering high volume but low linguistic diversity. While valuable for training, synthetic templates often fail to capture the subtle linguistic ambiguity that triggers safety-induced paralysis in deployment. In contrast, T3 prioritizes semantic depth over synthetic scale. By limiting our seed set to 454 expert-curated vignettes, we ensure each case targets a specific structural failure mode (e.g., distinguishing Confounding from Mediation) embedded in natural language. This design transforms the benchmark from a leaderboard metric into a diagnostic instrument that reports Utility and Safety separately, tracks calibrated abstention, and evaluates robustness under standardized pressure protocols.
@@
Prior work shows that preference optimization (RLHF) can induce sycophantic behavior, agreeing with a user’s mistaken belief to maximize approval (Sharma et al., 2023; Wei et al., 2023). This often stems from safety training objectives (Bai et al., 2022) and can lead to unfaithful reasoning traces (Turpin et al., 2023). Additionally, Lin et al. (2022) and the Inverse Scaling Prize (McKenzie et al., 2023) demonstrated that larger models can be less truthful or perform worse on tasks requiring disagreement. T3 studies these phenomena in the specific context of causal judgment: whether models endorse structurally invalid causal claims (Wolves) or refuse valid ones (Sheep), and how these behaviors vary across causal levels and under pressure.
@@ 2 Related Work
We position T3 at the intersection of causal reasoning benchmarks and sycophancy under preference tuning. While recent surveys highlight the potential of LLMs for causality (Kıcıman et al., 2023; Zhang et al., 2023; Chen et al., 2024), significant debate remains regarding whether these models possess genuine structural understanding or merely act as “causal parrots” (Zečević et al., 2023). Table 1 summarizes key differences from prior work.
@@ 2.1 Causal Reasoning Benchmarks
The evaluation of causal systems has long been grounded in structural learning principles (Spirtes et al., 2000; Schölkopf et al., 2021). Recent LLM benchmarks attempt to operationalize these principles. CLadder (Jin et al., 2023) generates queries from causal graphs, effectively testing the “do-calculus” (Pearl, 2000). However, real-world causal judgment often requires navigating informal ambiguity rather than formal symbols. This aligns with the “Causal Representation Learning” agenda (Schölkopf et al., 2021), which seeks to bridge high-level reasoning with low-level data representations.
@@
Evaluating models against the Causal Hierarchy Theorem (CHT) (Bareinboim et al., 2022) is a growing standard. CLadder (Jin et al., 2023) generates $\sim$10K queries from causal graphs across L1–L3. While its graph-first design provides formal control, synthetic variable construction yields puzzle-like scenarios distant from deployment contexts. CRASS (Frohberg and Binder, 2022) targets L3 counterfactuals but does not require explicit causal model construction. Recent work has also probed counterfactual consistency (Li et al., 2023) and logical modification (Huang et al., 2024), though often without the specific focus on failure-mode decomposition (Sheep vs. Wolf).
@@
The work of CORR2CAUSE (Jin et al., 2024) tests causal direction inference from correlations, but low performance suggests directionality from correlations alone is ill-posed for LLMs. At L1 (Association), the “Reversal Curse” (Berglund et al., 2023) demonstrates that models often fail to generalize $A\to B$ to $B\to A$, a fundamental associational deficit. Our work aligns with the “Epidemiology of LLMs” perspective (Plecko et al., 2025), which argues that models may memorize variable names but lack knowledge of the underlying observational distributions ($P(X,Y)$) required to identify traps like confounding (Rubin, 1974).
@@
e-CARE (Du et al., 2022) and BIG-Bench Causal Judgment (Srivastava et al., 2023) test narrative causal attribution. However, these often admit learned scripts rather than structural reasoning. T3 moves beyond narrative plausibility to test specific structural identifiability conditions (Pearl, 2009).
@@ 2.2 Sycophancy and Truthfulness
Sycophancy—where models agree with user biases to optimize for perceived helpfulness—is a known side-effect of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022). Sharma et al. (2023) and Turpin et al. (2023) demonstrated that models often rationalize incorrect answers if prompted with a biased context, even when using Chain-of-Thought reasoning (Wei et al., 2022). This phenomenon extends to multi-turn dialogues (Hong et al., 2025) and even objective domains like theorem proving (Petrov et al., 2025). Our work extends this to the causal domain, showing that models will endorse logical traps (like Simpson’s Paradox (Simpson, 1951)) if pressed.
@@
TruthfulQA (Lin et al., 2022) demonstrated that larger models can be less truthful due to mimicry of human misconceptions. The “Scaling Paradox” we observe at L3 mirrors the “Inverse Scaling Prize” findings (McKenzie et al., 2023), where larger models perform worse on tasks involving negation or counter-intuitive truths. T3 adds a new dimension to this literature: we identify ambiguity paralysis as a distinct mechanism of inverse scaling in safety-tuned models.

### v2

@@ 1.1 Relation to Prior Work
Two research threads inform our approach: formal causal assessment (spirtes2000causation; Schölkopf et al., 2021) and the study of alignment pathologies (bai2022const).
@@
Several benchmarks evaluate LLMs on Pearl’s hierarchy. CLadder (Jin and others, 2023) generates queries from causal graphs with oracle labels across rungs, offering high volume but low linguistic diversity. While valuable for training, synthetic templates often fail to capture the subtle linguistic ambiguity that triggers safety-induced paralysis in deployment. In contrast, T3 prioritizes semantic depth over synthetic scale. By limiting our seed set to 454 expert-curated vignettes, we ensure each case targets a specific structural failure mode (e.g., distinguishing Confounding from Mediation) embedded in natural language. This design transforms the benchmark from a leaderboard metric into a diagnostic instrument that reports Utility and Safety separately, tracks calibrated abstention, and evaluates robustness under standardized pressure protocols.
@@
Prior work shows that preference optimization (RLHF) can induce sycophantic behavior, agreeing with a user’s mistaken belief to maximize approval (Sharma and others, 2023; wei2023simple). This often stems from safety training objectives (bai2022const) and can lead to unfaithful reasoning traces (turpin2023language). Additionally, Lin et al. (2022) and the Inverse Scaling Prize (mckenzie2023inverse) demonstrated that larger models can be less truthful or perform worse on tasks requiring disagreement. T3 studies these phenomena in the specific context of causal judgment: whether models endorse structurally invalid causal claims (Wolves) or refuse valid ones (Sheep), and how these behaviors vary across causal levels and under pressure.
@@ 2 Related Work
We position T3 at the intersection of causal reasoning benchmarks and sycophancy under preference tuning. While recent surveys highlight the potential of LLMs for causality (Kıcıman et al., 2023; Zhang and others, 2023), significant debate remains regarding whether these models possess genuine structural understanding or merely act as “causal parrots” (Zečević et al., 2023). Table 1 summarizes key differences from prior work.
@@ 2.1 Causal Reasoning Benchmarks
The evaluation of causal systems has long been grounded in structural learning principles (spirtes2000causation; Schölkopf et al., 2021). Recent LLM benchmarks attempt to implement these principles. CLadder (Jin and others, 2023) generates queries from causal graphs, effectively testing the “do-calculus” (Pearl, 2000). However, real-world causal judgment often requires navigating informal ambiguity rather than formal symbols. This aligns with the “Causal Representation Learning” agenda (Schölkopf et al., 2021), which seeks to bridge high-level reasoning with low-level data representations.
@@
Evaluating models against the Causal Hierarchy Theorem (CHT) (Bareinboim et al., 2022) is a growing standard. CLadder (Jin and others, 2023) generates $\sim$10K queries from causal graphs across L1–L3. While its graph-first design provides formal control, synthetic variable construction yields puzzle-like scenarios distant from deployment contexts. CRASS (Frohberg and Binder, 2022) targets L3 counterfactuals but does not require explicit causal model construction. Recent work has also probed counterfactual consistency (Dehghanighobadi et al., 2025) and logical modification (Huang and others, 2024), though often without the specific focus on failure-mode decomposition (Sheep vs. Wolf).
@@
The work of CORR2CAUSE (Jin and others, 2024) tests causal direction inference from correlations, but low performance suggests directionality from correlations alone is ill-posed for LLMs. At L1 (Association), the “Reversal Curse” (Berglund and others, 2023) demonstrates that models often fail to generalize $A\to B$ to $B\to A$, a fundamental associational deficit. Our work aligns with the “Epidemiology of LLMs” perspective (Plecko et al., 2025), which argues that models may memorize variable names but lack knowledge of the underlying observational distributions ($P(X,Y)$) required to identify traps like confounding (Rubin, 1974).
@@
e-CARE (Du and others, 2022) and BIG-Bench Causal Judgment (Srivastava and others, 2023) test narrative causal attribution. However, these often admit learned scripts rather than structural reasoning. T3 moves beyond narrative plausibility to test specific structural identifiability conditions (Pearl, 2009).
@@ 2.2 Sycophancy and Truthfulness
Sycophancy—where models agree with user biases to optimize for perceived helpfulness—is a known side-effect of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; bai2022const). Sharma and others (2023) and turpin2023language demonstrated that models often rationalize incorrect answers if prompted with a biased context, even when using Chain-of-Thought reasoning (wei2022cot). This phenomenon extends to multi-turn dialogues (Hong et al., 2025) and even objective domains like theorem proving (Petrov et al., 2025). Our work extends this to the causal domain, showing that models will endorse logical traps (like Simpson’s Paradox (Simpson, 1951)) if pressed.
@@
TruthfulQA (Lin et al., 2022) demonstrated that larger models can be less truthful due to mimicry of human misconceptions. The “Scaling Paradox” we observe at L3 mirrors the “Inverse Scaling Prize” findings (mckenzie2023inverse), where larger models perform worse on tasks involving negation or counter-intuitive truths. T3 adds a new dimension to this literature: we identify ambiguity paralysis as a distinct mechanism of inverse scaling in safety-tuned models.